ANALYSIS OF THE F ASSST ROTATIONAL SPECTRUM OF S(CN)_{2}

\underline{Z. KISIEL}, O. DOROSH, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; I. R. MEDEVEDEV, M. BEHNKE, F. C. DE LUCIA, E. HERBST, M. WINNEWISSER, Department of Physics, The Ohio State University, Columbus, OH 43210.

We report a detailed analysis of the FASSST rotational spectrum of the S(CN)_{2} molecule, which has been recorded almost continuously from 110 to 375 GHz. S(CN)_{2} is of C_{2v} symmetry, is characterised by a low frequency bending mode of 120-130 cm^{−1}, and has four other normal modes with frequencies of less than 500 cm^{−1}. Only cm-wave transitions in the ground state of several isotopomers,^{a}^{,}^{b}^{,}^{c} and some transitions in the first excited bending state^{a} have previously been observed.

Rotational transitions in first excited states of all five lowest frequency normal have now been assigned, as well as in the bending progression up to v = 5, and in some combination states. The use of the efficient, newly written AABS package for Analysis and Assignment of Broadband Spectra allowed several thousand lines to be added also to the datasets for the ground states of the parent, and the ^{34}S isotopomers. The total number of lines measured in the FASSST spectrum of S(CN)_{2} is well over 10000. Octic level spectroscopic constants have been derived for the studied states, and their vibrational dependence will be discussed. Vibrational assignment was confirmed by means of ab initio inertial defect calculations.

aW.Arnold, H.Dreizler, H.D.Rudolph, Z. Naturforsch. 19a, 1428 (1964)

bL.Pierce, R.Nelson, R.Thomas, J. Chem. Phys. 43, 3423 (1965)

cH.M.Jemson, M.C.L.Gerry, J. Mol. Spectrosc. 124, 481 (1987)